OpenStudy (anonymous):

Newton currently has an account balance of $6,225.18. He opened the account 15 years ago with a deposit of $4,543.11. If the interest compounds daily, what is the interest rate on the account?

OpenStudy (anonymous):

what is the compounding formula you are given?

OpenStudy (anonymous):

The formula would be 6225.18=4543.11(1+r/365)^(365*15) I think. I just don't know how to solve it.

OpenStudy (anonymous):

looks good ...divide both sides by 4543.11 and exponentiate both sides with (1/(365*15))

OpenStudy (anonymous):

I'm still confused. 1.37=(1+r/365)^5475 What do I do from there?

OpenStudy (anonymous):

\[1.37^{1/5475}=((1+r/365)^{5475})^{1/5475}\]

OpenStudy (anonymous):

you good on this?

OpenStudy (anonymous):

Not really.

OpenStudy (anonymous):

what part?

OpenStudy (anonymous):

All of it. I can normally figure out problems like this, but when it comes to finding the rate in problems like this, I get stumped.

OpenStudy (anonymous):

I understand the beginning of the problem, but when you put the most recent equation, I got even more confused.

OpenStudy (anonymous):

\[(C ^{a})^{b} =C ^{ab} \]

OpenStudy (anonymous):

exponentiating both sides with (1/5475) gets rid of the exponent on the right side because 5475*(1/5475) =1

OpenStudy (anonymous):

So it would just be 1.37^5475=1+r/365 ?

OpenStudy (anonymous):

1.37^(1/5475) =1+r/365

OpenStudy (anonymous):

Ohh. Okay. So is the answer 364.02 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Omg. I got 1.00006=1+r/365 Then I multiplied 1.00006 by 365 and got 365.02 then subtracted 1 from both sides and got 364.02.

OpenStudy (anonymous):

hit these keys: 5 4 7 5 \[\frac{ 1 }{ x }\] memory store 1.37 \[y ^{x}\] memory recall - 1 * 365

OpenStudy (anonymous):

So 2.1% ?

OpenStudy (anonymous):

yep:)

OpenStudy (anonymous):

Yay. (: Finally.